There are 30 boys and 20 girls in a class. 20 boys and 10 girls use their own transport to reach the college. If a student is selected at random find the probability that the student selected is a boy or a student using his or her own transport to reach the college.
1
Expert's answer
2019-07-09T13:14:59-0400

P(A or B) = P(A) + P (B) - P(A and B)


P( x is a Boy) = 30 / 50 = 0.6

P(x is a driver) = (20 + 10) / 50 = 0.6

P(x is a Boy AND x is a driver) = 20 / 50 = 0.4


P(x is a Boy OR x is a driver) = 0.6 + 0.6 - 0.4 = 0.8


Answer: 0.8